| Operation Summary for February 3, 2012 - February 6, 2012 |
|
NuMI reported receiving 1.51 E18 protons on target for the period from 1/30/12 to 2/6/12. Friday February 3 The day shift began day four of a seven-day shutdown to replace a Main Injector (MI) quadrupole magnet (Q403). During the day shift, at 10:00 AM Interlock techs began safety system testing in Meson, they were finished at 10:42 AM. Operators established beam to the Booster dump at 3:51 PM. During the evening shift, Operations completed a search and secure of the Main Injector tunnel at 6:18 PM. Operators established beam to MiniBooNE at 8:29 PM. Linac RF Stations #2 (LRF2) tripped off at 10:00 PM, operators could not reset, experts were contacted and came in to investigate. Saturday February 4 The midnight shift began day five of the seven-day repair period. MiniBooNE was off due to LRF2 problem. During the midnight shift, LRF2 repair was completed at 12:24 AM, beam was then established to MiniBooNE. At 6:30 AM MiniBooNE was turned off in order to prepare the Main Injector for access to check on the status of the lambertson bakeout. LRF2 tripped off at 7:43 AM, operators could not reset, experts were contacted and came in to investigate. During the day shift, Main Injector access was completed at 8:05 AM. LRF2 repair was completed at 2:17 PM, beam was then re-established to MiniBooNE. During the evening shift, at 4:26 PM MiniBooNE was turned off in order to prepare the Main Injector for access to check on the status of the lambertson bakeout. Main Injector access was completed at 6:41 PM. Beam was then re-established to MiniBooNE. Sunday February 5 The midnight shift began day six of the seven-day repair period and with Operations monitoring beam to MiniBooNE. During the midnight shift, operations monitored beam to MiniBooNE. During the day shift, at 8:30 AM MiniBooNE was turned off in order to prepare the Main Injector for access to check on the status of the lambertson bakeout. Main Injector access was completed at 10:30 AM. Beam was then re-established to MiniBooNE. At 3:52 PM MiniBooNE was turned off once again to check on the lambertson bake. During the evening shift, Main Injector access finished at 5:14 PM. Beam was then re-established to MiniBooNE. Monday February 6 The midnight shift began day seven of the seven-day repair period and with Operations monitoring beam to MiniBooNE. During the midnight shift, Main Injector safety system testing began at 5:00 AM, thus halting beam to MiniBooNE. MI safety system testing was finished by 7:00 AM. Operations established beam to the Booster dump at 7:18 AM. Operations Coordinator (OpsCo) The OpsCo plan is to replace the PA for BRF18, end the MI lambertson bake, and let it cool off overnight, resume beam to MiniBooNE when possible. |
|
More Information Comments and Suggestions |
| last modified 02/06/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |